OpenStudy (anonymous):

Solve the equation by using factoring: x2 -14x + 49 =0 14 7 2 no solution

OpenStudy (anonymous):

just factor?

OpenStudy (anonymous):

\[x^2 -14x + 49 =0\]

OpenStudy (anonymous):

this factors pretty nicely

OpenStudy (anonymous):

what multiplies to +49 but adds up to -14?

OpenStudy (anonymous):

\[-7 \times -7 = +49\] \[-7 + - 7 = -14 \]

OpenStudy (anonymous):

\[x^2 -14x + 49 =0\] \[(x - 7) (x -7) = 0\] \[(x-7)^2 \] Now: \[x-7 = 0 \] SOLVE FOR X. x - 7 = 0 +7 +7 ------- ----- x + _____

OpenStudy (anonymous):

Sorry, typo. x = ____ not x + ____ LOL sorry :)

OpenStudy (anonymous):

7? @some_someone

OpenStudy (anonymous):

it factors to (x-7)(x-7) = 0 so yes, x = 7 they're repeated roots

OpenStudy (anonymous):

-7*-7 = 49 -7+(-7) = -14

OpenStudy (anonymous):

? you lost me